A quality control expert at Glotech computers wants to test their new monitors. The production manager claims they have a mean life of 8383 months with a variance of 100100.

If the claim is true, what is the probability that the mean monitor life would differ from the population mean by less than 1.71.7 months in a sample of 8989 monitors? Round your answer to four decimal places.
Answer #1

Given: Mean() 83 Noriance Co2 =100 Standard deviation () Vanonce Sample mean CX) Sample sizecn) Differ by mean by lessthn 17Hence awe get PZ <I.60) & p(z-160) = 0-9452 = 00S48 Now Tule PCa<z <b) P Z<6)-PZ <a) p-1-60z<I60)= P(z160-P(2G 60) O,O548 0-